DK4L:

Jimmy rode his bike miles to the train station. He rode the train into downtown. He then rode his bike miles to get to work. How many miles total does Jimmy ride his bike on his way to work?

ZombieLover:

Wheres the numbers?

DK4L:

Idk what happened to them

ZombieLover:

What are the numbers? i can't answer without them

DK4L:

I answered that one all ready. But I need help with this

ZombieLover:

oh ok

DK4L:

Felipe planted some seeds for a science project. One of his plants grew 6 1/2 inches in the first week, and then it grew 5 3/4 inches during the second week. How much did the plant grow during these two weeks? The answer should be written as a proper mixed number and should be simplified, if possible.

ZombieLover:

Um ask one of these @mxddi3 , @darkknight , maybe they can help

DK4L:

ok thnx

darkknight:

6 1/2 = 13/2 = 26/4 5 3/4 = 23/4 (23+26)/4 = 49/4 now write as a mixed number

DK4L:

12 1/4 right

darkknight:

yes good job, : )